how do you find it and the steps to solve it -7=-2(x-1)
1
Expert's answer
2014-06-02T04:21:32-0400
Reverse the equality in -7=-2(x-1) in order to isolate x to the left hand side.

-7=-2(x-1) is equivalent to -2(x-1)=-7.

Divide both sides a constant to simplify the equation. Divide both sides of -2(x-1)=-7 by -2.
Any nonzero number divided by itself is one. -2/(-2)=1: x-1=-7/(-2).

Simplify the sign of -7/(-2). The sign of -7/(-2) is positive, so -7/(-2)=7/2.

x-1=7/2.

Isolate terms with x to the left hand side.
Add 1 to both sides: x+(1-1)=7/2+1.


Look for two terms that sum to zero. 1-1=0:
x=7/2+1.


Evaluate 7/2+1=9/2.


Answer. x=9/2=4.5.